# Characterizing the support of semiclassical measures for higher-dimensional cat maps

### v1

@@ 6.3. Preliminary calculations
To prove (6.13), (6.14), and (6.15), we will bound either #$\#(S_{k}\cap\mathbb{Z}^{2})$ or #$\#(S_{k}\cap\mathbb{Z}^{2}\setminus 0)$.
2. We first show (6.13). Suppose $l,l^{\prime}\in S_{k}\cap\mathbb{Z}^{2}$. Then $|a_{-}(l-l^{\prime})|\leq|a_{-}(l+c)|+|a_{-}(l^{\prime}+c)|\leq 4\sqrt{kh}$. Thus, if $l\neq l^{\prime}$, from Lemma 6.3, $|a_{+}(l+c)-a_{+}(l^{\prime}+c)|=|a_{+}(l-l^{\prime})|\geq\frac{c_{0}}{4\sqrt{kh}}.$ Combining this with $|a_{+}(l+c)|,|a_{+}(l^{\prime}+c)|\leq 2\lambda^{q}\sqrt{kh}$, we know
(6.17) #$\#(S_{k}\cap\mathbb{Z}^{2})\leq\frac{16\lambda^{q}kh}{c_{0}}+1.$
@@
(6.18) $\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}\leq C\lambda^{q}h\sum_{k=1}^{\infty}ke^{-k}+\sum_{k=1}^{\infty}e^{-k}\leq C(\lambda^{q}h+1).$
@@
3. Now we prove (6.14). Assume that $|c|\ll\min\{1,\lambda^{-q}(h\log h^{-1})^{-\frac{1}{2}}\}$ and $h\log h^{-1}\ll\lambda^{-q}$. Note that
$S_{k}\subset\left\{z\in\mathbb{R}^{2}:|a_{+}(z)|\leq 2\lambda^{q}\sqrt{kh}+|c|,\ |a_{-}(z)|\leq 2\sqrt{kh}+|c|\right\}\eqqcolon\tilde{S}_{k}.$
We claim that #$\#(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$ for $k\leq 2\log h^{-1}$. Indeed, for $l,l^{\prime}\in\tilde{S}_{k}\cap\mathbb{Z}^{2}$, $|a_{-}(l-l^{\prime})|\leq 4\sqrt{kh}+2|c|$. Then if $l\neq l^{\prime}$, by Lemma 6.3, $|a_{+}(l)-a_{+}(l^{\prime})|\geq\frac{c_{0}}{4\sqrt{kh}+2|c|}$. Since $|a_{+}(l)|,|a_{+}(l^{\prime})|\leq 2\lambda^{q}\sqrt{kh}+|c|$, we see
#$\#(\tilde{S}_{k}\cap\mathbb{Z}^{2})\leq c_{0}^{-1}(2\lambda^{q}\sqrt{kh}+|c|)(4\sqrt{kh}+2|c|)+1\leq C(\lambda^{q}kh+|c|\lambda^{q}\sqrt{kh}+|c|^{2})+1.$
As $|c|\ll 1$, to show #$\#(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$, it suffices to prove $kh+|c|\sqrt{kh}\ll\lambda^{-q}$. Since $|c|\ll\lambda^{-q}(h\log h^{-1})^{-\frac{1}{2}}$ and $h\log h^{-1}\ll\lambda^{-q}$, for $k\leq 2\log h^{-1}$,
$kh+|c|\sqrt{kh}\leq 2h\log h^{-1}+|c|\sqrt{2h\log h^{-1}}\ll\lambda^{-q}.$
Thus, #$\#(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$ for $k\leq 2\log h^{-1}$.
Noting that $0\in\tilde{S}_{k}$, see #$\#(S_{k}\cap\mathbb{Z}^{2}\setminus 0)\leq\#(\tilde{S}_{k}\cap\mathbb{Z}^{2}\setminus 0)=0$ for $k\leq 2\log h^{-1}$. Therefore, using (6.17),
(6.19) $\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2}\setminus 0)\cdot e^{-k}\leq\sum_{2\log\frac{1}{h}<k}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}\leq C\lambda^{q}h\sum_{2\log\frac{1}{h}<k}e^{-k/2}+\sum_{2\log\frac{1}{h}<k}e^{-k}\leq C\lambda^{q}h.$
@@
[…]
@@
We apply (6.15) from Lemma 6.5 with $c=0$ to know
@@
[…]
@@
From these formulas, it is clear that $S_{1}(\lambda)\coloneqq\lim_{h\rightarrow 0}S_{1}(\lambda,\alpha,\varphi,P)=\sum_{t=-\infty}^{\infty}\frac{2}{\lambda^{t}+\lambda^{-t}}$ and $S_{2}(\lambda)\coloneqq\lim_{h\rightarrow 0}S_{2}(\lambda,\alpha,\varphi,P)$ exist and $S_{1}$ is strictly positive. ∎
@@
Figure 2. Plot of the concentration in position space of the eigenfunction defined in (6.5) corresponding to $A=\begin{bmatrix}5&2\\ 2&1\end{bmatrix}$, $k=6$, and $N_{k}=6930$. To create a more intelligible picture, we have shifted the eigenfunction to be centered at $(1/2,1/2)$, instead of the origin.
@@
$\lim_{h\rightarrow 0}\frac{\left\langle\left(U_{j/N}\otimes U_{k/N}\right)u,u\right\rangle_{\mathcal{H}}}{\left\langle u,u\right\rangle_{\mathcal{H}}}=\begin{cases}1&\text{ if }j=k=0\\ \frac{1}{2}&\text{ if }j=0,k\neq 0\\ \frac{1}{2}&\text{ if }j\neq 0,k=0\\ 0&\text{ if }j,k\neq 0.\end{cases}$
@@
First, recall the definition of $u_{I}$ from (6.20). If $I$ is an interval with $I\subset I_{0}$, then $|I|\leq 2\Theta(h)$. Then by Lemma 6.6, we know $\|u_{I}\|_{\mathcal{H}}=o(1)$. Therefore for all $j,k\in\mathbb{Z}^{2}$ (including $j=k=0$) and $I\subset I_{0}$,
@@
(6.24)$=\frac{1}{P}\sum_{s=-\frac{P}{2}}^{\frac{P}{2}-1}\sum_{t=-\frac{P}{4}}^{\frac{P}{4}-1}e^{-i\frac{\varphi}{P}(2t-2s)}\Big{\langle}G_{N},M_{N}^{s-t}G_{N}\Big{\rangle}_{\mathcal{H}}^{2}+o(1).$ ) .
@@
3. Finally, we examine the case where $j,k\neq 0$. Let $T_{*}\coloneqq[-\frac{P}{2},-\frac{P}{4}-1]\cup[\frac{P}{4},\frac{P}{2}-1]$. We note that $[-P/2,P/2-1]\bmod P=T_{*}\cup(T_{*}+P/2)\bmod P$. Using this decomposition to split up the sum over $t$ in (6.24) and exploiting the $P$-periodicity of $M_{N}$, we have
[…]
@@ 1. First note that it suffices to show that
By (6.4) and the fact that $U_{\omega}U_{\omega^{\prime}}=e^{\frac{i}{2h}\sigma(\omega,\omega^{\prime})}U_{\omega+\omega^{\prime}}$, we know
[…]
@@
2. We begin by proving (6.27). From Lemma 6.2, recalling that $R=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$,
$\left\langle U_{\omega}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}=\frac{\sqrt{2}}{\sqrt{\lambda^{q}+\lambda^{-q}}}e^{-\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}e^{i\frac{\left\langle A^{q}R\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}.$
@@
(6.29) $\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}-\left\langle G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|\leq\left|1-e^{-\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}+i\frac{\left\langle A^{q}R\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}\right|.$
Since $|1-e^{z}|\leq C|z|$ for $|z|\leq 1$, to further bound (6.29), we want to show $|\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}-i\frac{\left\langle A^{q}R\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}|=o(P^{-1})\leq 1$. In fact, we prove the stronger claim: for any fixed, $t$-independent matrix $B$ and for all $t\in I_{L}$, $\frac{|\left\langle A^{\pm q}B\omega,\omega\right\rangle|}{2h(\lambda^{q}+\lambda^{-q})}=o(P^{-1})$. We see this in the following calculation:
$\frac{|\left\langle A^{\pm q}B\omega,\omega\right\rangle|}{2h(\lambda^{q}+\lambda^{-q})}\leq\frac{|A^{\pm q}B\omega||\omega|}{2h\lambda^{|q|}}\leq\frac{C|A^{t}j/N|^{2}}{h}\leq\frac{C\lambda^{2|t|}}{N}\leq\frac{C\lambda^{\frac{P}{2}-2\Theta(h)}}{N}\leq C\lambda^{-2\Theta(h)}=o(P^{-1}).$
Thus setting $B=I$ and $B=R$, from (6.29) we have
$\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}-\left\langle G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|$ $\leq C\left|\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}-i\frac{\left\langle A^{q}R\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}\right|$
@@ 1. First note that it suffices to show
From (2.6), (6.4), and the fact that $U_{\omega}U_{\omega^{\prime}}=e^{\frac{i}{2h}\sigma(\omega,\omega^{\prime})}U_{\omega+\omega^{\prime}}$, we have
[…]
@@
(6.31)$=\frac{C}{P}\sum_{|q|\leq\frac{\Theta(h)}{2}}\sum_{t\in I_{E}}\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|+\frac{C}{P}\sum_{|q|=\frac{\Theta(h)}{2}+1}^{\frac{P}{2}}\sum_{t\in I_{E}}\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|.$
@@
Importantly, as $j\in\mathbb{Z}^{2}\setminus 0$, $a_{+}(j),a_{-}(j)\neq 0$. Thus, letting $C$ and $C^{\prime}$ be constants that may change from line to line, for $t\in I_{E}$:

### v2

@@ 6.3. Preliminary calculations
To prove (6.13), (6.14), and (6.15), we will bound either $(S_{k}\cap\mathbb{Z}^{2})$ or $(S_{k}\cap\mathbb{Z}^{2}\setminus 0)$.
2. We first show (6.13). Suppose $l,l^{\prime}\in S_{k}\cap\mathbb{Z}^{2}$. Then $|a_{-}(l-l^{\prime})|\leq|a_{-}(l+r)|+|a_{-}(l^{\prime}+r)|\leq 4\sqrt{kh}$. Thus, if $l\neq l^{\prime}$, from Lemma 6.3, $|a_{+}(l+r)-a_{+}(l^{\prime}+r)|=|a_{+}(l-l^{\prime})|\geq\frac{c_{0}}{4\sqrt{kh}}.$ Combining this with $|a_{+}(l+r)|,|a_{+}(l^{\prime}+r)|\leq 2\lambda^{q}\sqrt{kh}$ and $|a_{-}(l+r)|\leq 2\sqrt{kh}$, we know
(6.17) $(S_{k}\cap\mathbb{Z}^{2})\leq\left(2(2\lambda^{q}\sqrt{kh})\left(\frac{4\sqrt{kh}}{c_{0}}\right)+1\right)\cdot\left(2(2\sqrt{kh})+1\right)\leq C\left(k^{3/2}h\lambda^{q}+\sqrt{k}\right).$
@@
(6.18) $\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}\leq C\lambda^{q}h\sum_{k=1}^{\infty}k^{3/2}e^{-k}+\sum_{k=1}^{\infty}\sqrt{k}e^{-k}\leq C(\lambda^{q}h+1).$
@@
3. Now we prove (6.14). Assume that $|r|\ll\min\{1,\lambda^{-q}(h\log h^{-1})^{-\frac{1}{2}}\}$ and $h\log h^{-1}\ll\lambda^{-q}$. Note that
$S_{k}\subset\left\{z\in\mathbb{R}^{2}:|a_{+}(z)|\leq 2\lambda^{q}\sqrt{kh}+|r|,\ |a_{-}(z)|\leq 2\sqrt{kh}+|r|\right\}\eqqcolon\tilde{S}_{k}.$
We claim that $(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$ for $k\leq 2\log h^{-1}$. Indeed, for $l,l^{\prime}\in\tilde{S}_{k}\cap\mathbb{Z}^{2}$, $|a_{-}(l-l^{\prime})|\leq 4\sqrt{kh}+2|r|$. Then if $l\neq l^{\prime}$, by Lemma 6.3, $|a_{+}(l)-a_{+}(l^{\prime})|\geq\frac{c_{0}}{4\sqrt{kh}+2|r|}$. Since $|a_{+}(l)|,|a_{+}(l^{\prime})|\leq 2\lambda^{q}\sqrt{kh}+|r|$ and $|a_{+}(l)|\leq 2\sqrt{kh}+r\ll 1$, we see
$(\tilde{S}_{k}\cap\mathbb{Z}^{2})\leq 2\frac{4\sqrt{kh}+2|r|}{c_{0}}(2\lambda^{q}\sqrt{kh}+|r|)+1\leq C(\lambda^{q}kh+|r|\lambda^{q}\sqrt{kh}+|r|^{2})+1.$
As $|r|\ll 1$, to show $(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$, it suffices to prove $kh+|r|\sqrt{kh}\ll\lambda^{-q}$. Since $|r|\ll\lambda^{-q}(h\log h^{-1})^{-\frac{1}{2}}$ and $h\log h^{-1}\ll\lambda^{-q}$, for $k\leq 2\log h^{-1}$,
$kh+|r|\sqrt{kh}\leq 2h\log h^{-1}+|r|\sqrt{2h\log h^{-1}}\ll\lambda^{-q}.$
Thus, $(\tilde{S}_{k}\cap\mathbb{Z}^{2})<2$ for $k\leq 2\log h^{-1}$.
Noting that $0\in\tilde{S}_{k}$, see $(S_{k}\cap\mathbb{Z}^{2}\setminus 0)\leq\#(\tilde{S}_{k}\cap\mathbb{Z}^{2}\setminus 0)=0$ for $k\leq 2\log h^{-1}$. Therefore, using (6.17),
(6.19) $\sum_{k=1}^{\infty}\#(S_{k}\cap\mathbb{Z}^{2}\setminus 0)\cdot e^{-k}\leq\sum_{2\log\frac{1}{h}<k}\#(S_{k}\cap\mathbb{Z}^{2})\cdot e^{-k}\leq C\lambda^{q}h\sum_{2\log\frac{1}{h}<k}e^{-k/2}+\sum_{2\log\frac{1}{h}<k}e^{-k/2}\leq C\lambda^{q}h.$
@@
4. Finally, we show (6.15). Assume $|r|\ll 1$. If $1\leq\lambda^{-q}(h\log h^{-1})^{-1/2}$, then $|r|\ll\min\{1,\lambda^{-q}(h\log h^{-1})^{-\frac{1}{2}}\}$ and $h\log h^{-1}\ll\lambda^{-q}$. Thus we have (6.14), a stronger statement than (6.15).
@@
We apply (6.13) and (6.15) from Lemma 6.5 with $r=0$ to know
@@
$\leq\frac{C}{P}\left(h\frac{\lambda^{\frac{\alpha+1}{2}}-\alpha(\lambda^{\frac{1}{2}}-1)-\lambda^{\frac{1}{2}}}{(\lambda^{\frac{1}{2}}-1)^{2}}+(hP)^{\frac{1}{4}}\alpha^{2}\right)$
@@
(6.21) $=\frac{1}{P}\sum_{t=-\alpha+1}^{\alpha-1}\left(\alpha-|t|\right)e^{-\frac{i\varphi}{P}2t}\left\langle M^{t}_{A}G_{h},G_{h}\right\rangle_{L^{2}}^{2}+o(1)$
@@
Recall that $\varphi\in[-\pi,\pi)$; in particular, $\varphi/P$ converges to 0. From these formulas, it is clear that $S_{1}(\lambda)\coloneqq\lim_{h\rightarrow 0}S_{1}(\lambda,\alpha,\varphi,P)=\sum_{t=-\infty}^{\infty}\frac{2}{\lambda^{t}+\lambda^{-t}}$ and $S_{2}(\lambda)\coloneqq\lim_{h\rightarrow 0}S_{2}(\lambda,\alpha,\varphi,P)$ exist and $S_{1}$ is strictly positive. ∎
@@
Figure 3. Plot of the concentration in position space of the eigenfunction defined in (6.5) corresponding to $A=\begin{bmatrix}5&2\\ 2&1\end{bmatrix}$, $k=6$, and $N_{k}=6930$. Specifically, we blurred the squared absolute value of the eigenfunction. This blurring corresponds to taking the macroscopic limit, as we can only integrate against slowly varying functions. To create a more intelligible picture, we have shifted the eigenfunction to be centered at $(1/2,1/2)$, instead of the origin.
@@
[…]
@@
First, recall the definition of $u_{I}$ from (6.20). If $I$ is an interval with $I\subset I_{0}$, then $|I|\leq 2\Theta(h)$. Then by Lemma 6.6 and (6.22), we know $\|u_{I}\|_{\mathcal{H}}=o(1)$. Therefore for all $j,k\in\mathbb{Z}^{2}$ (including $j=k=0$) and $I\subset I_{0}$,
@@
[…]
@@
3. Finally, we examine the case where $j,k\neq 0$. Let $T_{*}\coloneqq[-\frac{P}{2},-\frac{P}{4}-1]\cup[\frac{P}{4},\frac{P}{2}-1]$. We note that $P/2,P/2-1]\bmod P=T_{*}\cup(T_{*}+P/2)\bmod P$. Using this decomposition to split up the sum over $t$ in (6.24) and exploiting the $P$-periodicity of $M_{N}$, we have
(6.24) $=\frac{1}{P}\sum_{s=-\frac{P}{2}}^{\frac{P}{2}-1}\sum_{t\in T_{*}}e^{-i\frac{\varphi}{P}(2t-2s)}\Big\langle U_{j/N}M^{t}_{N}G_{N},M_{N}^{s}G_{N}\Big\rangle_{\mathcal{H}}\Big\langle U_{k/N}M^{t+\frac{P}{2}}_{N}G_{N},M_{N}^{s+\frac{P}{2}}G_{N}\Big\rangle_{\mathcal{H}}$
@@ 1. First note that it suffices to show that
By (6.4),  (2.11),  (2.13), and the fact that $U_{\omega}U_{\omega^{\prime}}=e^{\frac{i}{2h}\sigma(\omega,\omega^{\prime})}U_{\omega+\omega^{\prime}}$, we know
(6.26) $\leq\frac{C}{P}\sum_{|s|\leq\frac{P}{2}}\sum_{t\in I_{L}}\left|\left\langle U_{A^{-t}j/N}G_{N},M_{N}^{q_{s,t}}G_{N}\right\rangle_{\mathcal{H}}-\left\langle G_{N},M^{q_{s,t}}_{N}G_{N}\right\rangle_{\mathcal{H}}\right|$
@@
2. We begin by proving (6.27). From Lemma 6.2, recalling that $J=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$,
$\left\langle U_{\omega}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}=\frac{\sqrt{2}}{\sqrt{\lambda^{q}+\lambda^{-q}}}e^{-\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}e^{i\frac{\left\langle A^{q}J\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}.$
@@
(6.29) $\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}-\left\langle G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|\leq\left|1-e^{-\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}+i\frac{\left\langle A^{q}J\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}}\right|.$
Since $|1-e^{z}|\leq C|z|$ for $|z|\leq 1$, to further bound (6.29), we want to show $|\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}-i\frac{\left\langle A^{q}J\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}|=o(P^{-1})\leq 1$. In fact, we prove the stronger claim: for any fixed, $t$-independent matrix $B$ and for all $t\in I_{L}$, $\frac{|\left\langle A^{\pm q}B\omega,\omega\right\rangle|}{2h(\lambda^{q}+\lambda^{-q})}=o(P^{-1})$. We see this in the following calculation:
$\frac{|\left\langle A^{\pm q}B\omega,\omega\right\rangle|}{2h(\lambda^{q}+\lambda^{-q})}\leq\frac{|A^{\pm q}B\omega||\omega|}{2h\lambda^{|q|}}\leq\frac{C|A^{-t}j/N|^{2}}{h}\leq\frac{C\lambda^{2|t|}}{N}\leq\frac{C\lambda^{\frac{P}{2}-2\Theta(h)}}{N}\leq C\lambda^{-2\Theta(h)}=o(P^{-1}).$
Thus setting $B=I$ and $B=J$, from (6.29) we have
$\left|\left\langle U_{A^{-t}j/N}G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}-\left\langle G_{h},M_{A}^{q}G_{h}\right\rangle_{L^{2}}\right|$ $\leq C\left|\frac{\left\langle A^{-q}\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}-i\frac{\left\langle A^{q}J\omega,\omega\right\rangle}{2h(\lambda^{q}+\lambda^{-q})}\right|$
@@ 1. First note that it suffices to show
From (2.6), (6.4), and the facts that $U_{\omega}U_{\omega^{\prime}}=e^{\frac{i}{2h}\sigma(\omega,\omega^{\prime})}U_{\omega+\omega^{\prime}}$ and $\|G_{N}\|_{\mathcal{H}}\leq C$ uniformly in $N$, we have
(6.30) $=\frac{1}{P}\sum_{|s|\leq\frac{P}{2}}\sum_{t\in I_{E}}\left|\Big\langle U_{j/N}M_{N}^{t}G_{N},M_{N}^{s}G_{N}\Big\rangle_{\mathcal{H}}\Big\langle U_{k/N}M_{N}^{t+\frac{P}{2}}G_{N},M_{N}^{s+\frac{P}{2}}G_{N}\Big\rangle_{\mathcal{H}}\right|$
@@
[…]
@@
Importantly, as $j\in\mathbb{Z}^{2}\setminus 0$ and the slopes of $v_{+},v_{-}$ are irrational, we know $a_{+}(j),a_{-}(j)\neq 0$. Thus, letting $C$ and $C^{\prime}$ be constants that may change from line to line, for $t\in I_{E}$: